OpenStudy (anonymous):

(k+1/k)^2=16, then k^2+1/k^2

OpenStudy (campbell_st):

the expansion of (k + 1/k)^2 = 16 is k^2 + 2 + 1/k^2 = 16 you should be able to find k^2 + 1/k^2 now...

OpenStudy (anonymous):

i got k^2+1/k^2=14

OpenStudy (campbell_st):

correct.... well done